OpenStudy (anonymous):

what is the product of the first row,second column [-3 6] x [2 3 ] = [ ] 2-4 6 9

myininaya (myininaya):

\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] \times \left[\begin{matrix}u & v \\ w & x\end{matrix}\right]=\left[\begin{matrix}au+bw & av+bx \\ cu+dw & cv+dx\end{matrix}\right]\]